OpenStudy (marianabelli):

CD is perpendicular to AB and passes through point C(5, 12). If the coordinates of A and B are (-10, -3) and (7, 14), respectively, the x- intercept CD of is . The point lies on CD .

OpenStudy (marianabelli):

Options: 1 blank a.(12,0) b.(15,0) c.(17,0) d.(19,0) 2 blank. a.(-5,24) b.(-2,19) c.(7,-10) d.(8,11)

OpenStudy (marianabelli):

Please help

OpenStudy (qwertty123):

@hartnn

OpenStudy (sshayer):

slope of AB\[=\frac{ 14-(-3) }{ 7-(-10) }=\frac{ 17 }{ 17 }=1\] slope of line perpendicular to AB=-1/1=-1 eq. of CD is\[y-12=-1(x-5)\] y-12=-x+5 x+y=5+12=17 x+y=17 now you can check by substituting each point one by one and see which satisfies this eq.

OpenStudy (sshayer):

can you show me the work?

OpenStudy (sshayer):

are you here?